OpenStudy (chris215):

Find an exact value. cos 15°

OpenStudy (chris215):

is the answer (sqrt 6 + sqrt 2)/4

zepdrix (zepdrix):

Hmm

zepdrix (zepdrix):

\[\Large\rm \cos15=\cos\left(\frac{30}{2}\right)=\sqrt{\frac{1+\cos30}{2}}\]\[\Large\rm=\sqrt{\frac{1+\sqrt3/2}{2}}=\sqrt{\frac{2+\sqrt3}{4}}=\frac{\sqrt{2+\sqrt3}}{2}\]

zepdrix (zepdrix):

I'm a little confused where you got that 4 from in the bottom, Hmm

OpenStudy (chris215):

that answer is non of my choices though

zepdrix (zepdrix):

What are the choices? :D Maybe you can list them.

zepdrix (zepdrix):

These problems can be tough for multiple choice because it's hard to simplify things down to where they want.

OpenStudy (chris215):

yeah im pretty confused

zepdrix (zepdrix):

Oh ok.. so notice that I have a root within a root. So half angle formula was definitely a bad choice. There IS A WAY to denest roots, but it's definitely annoying. Let's go to our angle difference formula for cosine instead.

zepdrix (zepdrix):

\[\Large\rm \cos15=\cos(45-30)\]Ya?

zepdrix (zepdrix):

\[\Large\rm =\cos45\cos30+\sin45\sin30\]

OpenStudy (chris215):

yeah that's what i did

zepdrix (zepdrix):

\[\Large\rm =\frac{\sqrt2}{2}\cdot\frac{\sqrt3}{2}+\frac{\sqrt2}{2}\cdot\frac{1}{2}\]Oh ok good :)

zepdrix (zepdrix):

Ahh now I see where you got the 4 from ^^ hehe

zepdrix (zepdrix):

Ah yes your answer looks correct \c:/

zepdrix (zepdrix):

Sorry to ramble on so long lol

OpenStudy (chris215):

ohh ok but thanks so much I wasn't sure if I did it right or not